OpenStudy (anonymous):

Determine the standard form of the equation of the line that passes through (2, 7) and (-6, 0).

OpenStudy (anonymous):

Hold on. I'm writing it down.

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

Gee, if I had the slope form equation I could solve this but I'm not sure......is there any more to this problem?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

OK, do you know how to put this into slope format?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

OK, what is it? And if you don't know I'm looking in my math book. :)

OpenStudy (anonymous):

(y2 - y1) / (x2 - x1) (0-7) / (6 - 2)

OpenStudy (anonymous):

that was the first step, hopefully when were done you will know how to do this. :)

OpenStudy (anonymous):

-7 / 4 = Slope

OpenStudy (anonymous):

\[y-y1=\frac{ y2-y1 }{ x2-x1 }\left( x-x1 \right)\] (x1,y1)=(2,7) ( x2,y2 )=( -6,0 ) now you can try

OpenStudy (anonymous):

I'm working on this step by step so bear with me. :)

OpenStudy (anonymous):

what is B (where does this cross over the Y axis?)

OpenStudy (anonymous):

idk im lost now

OpenStudy (anonymous):

OK, when we draw the line between these 2 points on a graph where does the Line cross the Y axis, the vertical line?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

Do you have a graph? if so draw the two points and then tell me what NUMBER on the Y AXIS the line crosses through.

OpenStudy (anonymous):

I dont have a graph

OpenStudy (anonymous):

I'll attach it.

OpenStudy (anonymous):

It crosses the Y axis at the 5

OpenStudy (anonymous):

So with the slope form equation of Y = mx + b, we now have Y = -7/4 + 5

OpenStudy (anonymous):

Get it?

OpenStudy (anonymous):

yeah i got the slope intercept form part of that i dont get how to get standard form

OpenStudy (anonymous):

OK, I'll show you how to change them. )

OpenStudy (anonymous):

Now we since Y is a positive we'll subtract it from both sides: y =-7/4 + 5 ---- ----- -y -y

OpenStudy (anonymous):

Remember that we can NOT subtract y from 5 because they are NOT like terms! Also remember that Y is a negative! This will give us: -7/4 = -y + 5

OpenStudy (anonymous):

Follow so far?

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

OK either of you two other guys want to help?? Apparently I'm not a good teacher. :)

OpenStudy (anonymous):

no im just dumb lol

OpenStudy (anonymous):

That's Ok, I just did this lesson last night and my Dad, Math teacher, had to help me. Let me guess FLVS?

OpenStudy (anonymous):

Let me ask a friend for help, cause I'll be honest I'm kinda lost right now :)

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

thats fine thank you

OpenStudy (anonymous):

your welcome, I don't want to tell you the wrong thing.

OpenStudy (anonymous):

@aaronq

OpenStudy (anonymous):

let's hope one of them answers :)

OpenStudy (anonymous):

hopefully!

OpenStudy (anonymous):

Helpme100, I see you're here, want to help?

OpenStudy (anonymous):

Hey this might help, http://answers.yahoo.com/question/index?qid=20090302235850AA97pje

OpenStudy (anonymous):

that might help but I gotta go! Good Lusk!!! ;)